p(a) = .59 p(b) = .30 p(anb) =.21 find p(a'ub') and find p(a'nb')

p(a) = .59
p(b) = .30
p(anb) =.21

find p(a'ub') and find p(a'nb')
Answer #1

Given that - pca) = 0.59 P(b) = 0. 30 - PlanB) = 0.21 (anb) Plaub) = P(a) + P (6) - Prand) -= 0.59 +0.30-0.21 P Caub) - 0.68